OpenStudy (anonymous):

an ant is on a record player starting from one rest with a=0.4 rad/squared. what is the angular speed after 2 sec? how far has it travelled in 2sec? how long does it take to get around once?

OpenStudy (anonymous):

angular speed = initial speed + (angular acceleration x time) therefore, angular speed after 2 sec = 0 + (.4 x 2) = 0.8 rad/s Distance = 0.5 x angular acceleration x time^2 therefore, distance = 0.5 x .4 x 2 x 2 = .8 rad For one revolution distance = 360 degrees = 6.28318531radians therefore, 6.28 = 0.5 x .4 x t^2 t = sqrt (62.8/2) = sqrt (31.4) = 5.60 s

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

you are welcome! which country, standard are you from? Is this for homework?

OpenStudy (anonymous):

yeah one question I couldn't get

OpenStudy (anonymous):

does openstudy work well for you? I asked some question but can't get a response